# LLM in a flash: Efficient Large Language Model Inference with Limited Memory

### v2

@@ Abstract
Large language models (LLMs) are central to modern natural language processing, delivering exceptional performance in various tasks. However, their substantial computational and memory requirements present challenges, especially for devices with limited DRAM capacity. This paper tackles the challenge of efficiently running LLMs that exceed the available DRAM capacity by storing the model parameters in flash memory, but bringing them on demand to DRAM. Our method involves constructing an inference cost model that takes into account the characteristics of flash memory, guiding us to optimize in two critical areas: reducing the volume of data transferred from flash and reading data in larger, more contiguous chunks. Within this hardware-informed framework, we introduce two principal techniques. First, “windowing” strategically reduces data transfer by reusing previously activated neurons, and second, “row-column bundling”, tailored to the sequential data access strengths of flash memory, increases the size of data chunks read from flash memory. These methods collectively enable running models up to twice the size of the available DRAM, with a 4-5x and 20-25x increase in inference speed compared to naive loading approaches in CPU and GPU, respectively. Our integration of sparsity awareness, context-adaptive loading, and a hardware-oriented design paves the way for effective inference of LLMs on devices with limited memory.
@@ 1 Introduction
In recent years, large language models (LLMs), such as GPT-3 (Brown et al., 2020), OPT (Zhang et al., 2022b), and PaLM (Chowdhery et al., 2022), have demonstrated strong performance across a wide range of natural language tasks. However, the unprecedented capabilities of these models come with substantial computational and memory requirements for inference. LLMs can contain hundreds of billions or even trillions of parameters, which makes them challenging to load and run efficiently, especially on resource-constrained devices.
Currently, the standard approach is to load the entire model into DRAM (Dynamic Random Access Memory) for inference (Rajbhandari et al., 2021; Aminabadi et al., 2022). However, this severely limits the maximum model size that can be run. For example, a 7 billion parameter model requires over 14GB of memory just to load the parameters in half-precision floating point format, exceeding the capabilities of most edge devices.
Figure 1: Inference latency of 1 token when half the memory of the model is available. Our method selectively loads parameters on demand per token generation step. The latency is the time needed to load from flash multiple times back and forth during the generation of all tokens and the time needed for the computations, averaged over all generated tokens.
@@
To address this limitation, we propose to store the model parameters in flash memory, which is at least an order of magnitude larger than DRAM. Then, during inference, we directly load the required subset of parameters from the flash memory, avoiding the need to fit the entire model in DRAM. Our method is built on the top of recent works that have shown LLMs exhibit a high degree of sparsity in the Feed Forward Network (FFN) layers, with models like OPT (Zhang et al., 2022b), Falcon (Almazrouei et al., 2023), and Persimmon (Elsen et al., 2023), exhibiting more than 90% sparsity (Mirzadeh et al., 2023; Liu et al., 2023b). We exploit this sparsity to selectively load only parameters from flash memory that either have non-zero input or are predicted to have non-zero output. Specifically, we discuss a hardware-inspired cost model that includes flash memory, DRAM, and compute (CPU or GPU). Then, we introduce two complementary techniques to minimize data transfer and maximize flash memory throughput:
  • Windowing: We load and temporarily cache parameters for only the past few tokens, reusing aggregate sparsity structure predicted over the past few tokens. This sliding window approach reduces the number of IO requests to load weights.
  • Row-column bundling: We store a concatenated row and column of the up-projection and down-projection layers to read bigger contiguous chunks from flash memory. This increases throughput by reading larger chunks.
To further minimize the number of weights to be transferred from flash memory to DRAM, we also employ methods to predict FFN sparsity and avoid loading zeroed-out parameters, akin to approaches documented in Deja Vu (Li and Lu, 2023). Together, windowing and sparsity prediction allow us to load only  2% of the FFN layer from flash for each inference query. We also propose a static memory preallocation to minimize transfers within DRAM and reduce inference latency. Our load from flash cost model captures the tradeoff between loading less data and reading bigger chunks. Optimizing this cost model and selectively loading parameters on demand yields flash loading strategies that can run models 2x larger than the device’s DRAM capacity and speed up inference by 4-5x and 20-25x compared to naive implementation in CPU and GPU, respectively. It significantly outperforms the baseline approach, which reloads the model’s weights on every forward pass.
@@ 2 Flash Memory & LLM Inference
In this section, we explore the characteristics of memory storage systems (e.g., flash, DRAM), and their implications for large language model (LLM) inference. Our aim is to elucidate the challenges and hardware-specific considerations essential for algorithm design, particularly in optimizing inference when working with flash memory.
@@ 2.2 Read Throughput
Flash memory systems perform optimally with large sequential reads. For instance, benchmarks on an Apple MacBook Pro M2 with 2TB flash demonstrate speeds exceeding 6GiB/s for a 1GiB linear read of an uncached file. However, this high bandwidth is not replicated for smaller, random reads due to the inherent multi-phase nature of these reads, encompassing the operating system, drivers, interrupt handling, and the flash controller, among others. Each phase introduces latency, disproportionately affecting smaller reads.
To circumvent these limitations, we advocate two primary strategies, which can be employed jointly. The first involves reading larger chunks of data. For smaller blocks, a substantial part of the overall read time is spent waiting for data transfer to begin. This is often referred to as latency to first byte. This latency reduces the overall throughput of each read operation considerably, because the overall measured throughput has to take into account not just the speed of transfer once it begins, but the latency before it begins as well, which penalizes small reads. This means that if we coalesce the reads for rows and colums of the FFN matrices, we can pay the latency cost only once for any given row/column pair in both matrices, and higher throughput can be realized. This principle is depicted in Figure 1(b). Perhaps a counterintuitive yet interesting observation is that in some scenarios, it will be worthwhile to read more than needed (but in larger chunks) and then discard, than only reading strictly the necessary parts but in smaller chunks. The second strategy leverages parallelized reads, utilizing the inherent parallelism within storage stacks and flash controllers. Our results indicate that throughputs appropriate for sparse LLM inference are achievable on modern off-the-shelf hardware using 32KiB or larger random reads across multiple threads.
Motivated by the challenges described in this section, in section 3, we propose methods to optimize data transfer volume and enhance read throughput to significantly enhance inference speeds.
@@
Figure 3: (a) Preactivations of tokens in one sequence in OPT 6.7B. The blue graph shows preactivation of elements that predictor detected positive while the green graph is for up projection. As it can be seen most of the False Positives are close to 0 and False Negatives constitute a small portion of the elements. (b) A small low rank predictor finds out which intermediate neurons are going to be activated instead of running heavy up projection.
@@ 3 Load From Flash
Our proposed solutions for reducing latency under memory constraints are categorized into three strategic areas, each targeting a specific aspect of the latency:
  • Reducing Data Load: Aiming to decrease latency associated with flash I/O operations by loading less data [^1] It is notable that, by data we mean weights of the neural network. However, our developed techniques can be easily generalized to other data types transferred and used for LLM inference, such as activations or KV cache, as suggested by Sheng et al. (2023)..
@@
It is important to note that our focus is not on the compute aspect of the process, as it is orthogonal to the core concerns of our work. This delineation allows us to concentrate on optimizing flash memory interactions and memory management to achieve efficient inference on memory-constrained devices.
Finally, we will elaborate on the implementation of these strategies in subsequent sections.
@@ 3.1 Reducing Data Transfer
Our methodology leverages the inherent activation sparsity found in Feed-Forward Network (FFN) models, as documented in preceding research. The OPT 6.7B model, for instance, exhibits a notable 97% sparsity within its FFN layer. Similarly, the Falcon 7B model has been adapted through fine-tuning, which involves swapping their activation functions to ReLU, resulting in 95% sparsity while being almost similar in accuracy (Mirzadeh et al., 2023). In light of this information, our approach involves the iterative transfer of only the essential, non-sparse data from flash memory to DRAM for processing during inference.
While we employ the 7B models as practical examples to elucidate our approach, our findings are adaptable, and they can be extrapolated to both larger and smaller scale models.
Selective Persistence Strategy. We opt to retain the embeddings and matrices within the attention mechanism of the transformer constantly in RAM. For the Feed-Forward Network (FFN) portions, only the non-sparse segments are dynamically loaded into DRAM as needed. Storing attention weights, which constitute approximately one-third of the model’s size, in memory, allows for more efficient computation and quicker access, thereby enhancing inference performance without the need for full model loading.
Anticipating ReLU Sparsity. The ReLU activation function naturally induces over 90% sparsity in the FFN’s intermediate outputs, which reduces the memory footprint for subsequent layers that utilize these sparse outputs. However, the preceding layer, namely the up project for OPT and Falcon, must be fully present in memory. To avoid loading the entire up projection matrix, we follow  Liu et al. (2023b), and employ a low-rank predictor to identify the elements zeroed by ReLU (see Figure 2(b)). In contrast to their work, our predictor needs only the output of the current layer’s attention module, and not the previous layer’s FFN module. We have observed that postponing the prediction to current layer is sufficient for hardware aware weight loading algorithm design, but leads to more accurate outcome due to deferred inputs. We thereby only load elements indicated by the predictor.
Table 1: Using predictors doesn’t change the accuracy of zero-shot metrics significantly as predictor of each layer accurately identifies sparsity
@@
Figure 4: (a) Aggregated neuron use of the tenth layer of Falcon 7B. As it can be seen the slope of aggregated neuron use is decreasing. Other layers exhibit the same pattern. (b) Instead of deleting neurons that brought to DRAM we keep the active neurons of past 5 tokens: when the new token "Was" is being processed only a few amount of data needs to be changed.
Neuron Data Management via Sliding Window Technique. In our study, we define an active neuron as one that yields a positive output in our low rank predictor model. Our approach focuses on managing neuron data by employing a Sliding Window Technique. This technique entails maintaining a DRAM cache of of only the weight rows that were predicted to be required by the the recent subset of input tokens. The key aspect of this technique is the incremental loading of neuron data that differs between the current input token and its immediate predecessors. This strategy allows for efficient memory utilization, as it frees up memory resources previously allocated to cached weights required by tokens that are no longer within the sliding window (as depicted in Figure 3(b)).
From a mathematical standpoint, let $s_{\text{agg}}(k)$ denote the cumulative use of neuron data across a sequence of $k$ input tokens. Our memory architecture is designed to store an average of $s_{\text{agg}}(k)$ in DRAM. As we process each new token, the incremental neuron data, which is mathematically represented as $s_{\text{agg}}(k+1)-s_{\text{agg}}(k)$, is loaded from flash memory into DRAM. This practice is grounded in the observed trend of decreasing aggregated neuron usage over time. Consequently, larger values of $k$ result in a lesser volume of data being loaded for each new token. (refer to Figure 3(a)), while smaller values of $k$ can help conserve DRAM that is used to store the cached weights. In determining the size of the sliding window, the aim is to maximize it within the constraints imposed by the available memory capacity.
3.2 Improving Transfer Throughput with Increased Chunk Sizes
@@
Bundling Columns and Rows. For OPT and Falcon models, the usage of the $i$th column from the upward projection and the $i$th row from the downward projection coincides with the activation of the $i$th intermediate neuron. Consequently, by storing these corresponding columns and rows together in flash memory, we can consolidate the data into larger chunks for reading. Refer to Figure 5 for an illustration of this bundling approach. If each element of weights of the network is stored in num_bytes such bundling doubles the chunk size from $d_{model}\times$num_bytes to $2d_{model}\times$num_bytes as shown in Figure 5. Our analysis and experiment show this increases the throughput of the model.

### v3

@@ Abstract
Large language models (LLMs) are central to modern natural language processing, delivering exceptional performance in various tasks. However, their substantial computational and memory requirements present challenges, especially for devices with limited DRAM capacity. This paper tackles the challenge of efficiently running LLMs that exceed the available DRAM capacity by storing the model parameters in flash memory, but bringing them on demand to DRAM. Our method involves constructing an inference cost model that takes into account the characteristics of flash memory, guiding us to optimize in two critical areas: reducing the volume of data transferred from flash and reading data in larger, more contiguous chunks. Within this hardware-informed framework, we introduce two principal techniques. First, “windowing” strategically reduces data transfer by reusing previously activated neurons, and second, “row-column bundling”, tailored to the sequential data access strengths of flash memory, increases the size of data chunks read from flash memory. These methods collectively enable running models up to twice the size of the available DRAM, with up to 4x and 20x increase in inference speed compared to naive loading approaches in CPU and GPU, respectively. Our integration of sparsity awareness, context-adaptive loading, and a hardware-oriented design paves the way for effective inference of LLMs on devices with limited memory.
@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated strong performance across a wide range of natural language tasks (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023a; Jiang et al., 2023; Gemini Team, 2023). However, the unprecedented capabilities of these models come with substantial computational and memory requirements for inference. LLMs can contain hundreds of billions or even trillions of parameters, which makes them challenging to load and run efficiently, especially on personal devices.
Currently, the standard approach is to load the entire model into DRAM (Dynamic Random Access Memory) for inference (Rajbhandari et al., 2021; Aminabadi et al., 2022). However, this severely limits the maximum model size that can be run. For example, a 7 billion parameter model requires over 14GB of memory just to load the parameters in half-precision floating point format, exceeding the capabilities of most personal devices such as smartphones. While it is possible to employ techniques such as quantization to reduce the model size, still, this cannot address the main limitation of loading the entire model into DRAM.
Figure 1: Average inference latency for a single token when only half of the model’s memory is available: Our method selectively loads parameters on demand for each token generation step. The latency represents the time required to repeatedly load parameters from flash memory, combined with the time needed for computations.
@@
To address this limitation, we propose to store the model parameters in flash memory, which is at least an order of magnitude larger than DRAM. Then, during inference, we directly load the required subset of parameters from the flash memory, avoiding the need to fit the entire model in DRAM. To this end, our work makes several contributions:
  • First, we study the hardware characteristics of storage systems (e.g., flash, DRAM). We show that hardware constraints such as capacity and bandwidth limitations can have significant considerations when designing efficient algorithms for serving LLMs from flash (Section  2).
  • Motivated by our findings, we propose several techniques that can help with (i) reducing the required data transfer, (ii) increasing the transfer throughput, and (iii) managing loaded parameters efficiently in DRAM (Section  3).
  • Finally, as partially demonstrated in Figure  1, we show that our proposed techniques for optimizing the cost model and selectively loading parameters on demand allows us to run models 2x larger than the device’s DRAM capacity and speed up inference up to 4x, 7x, and 20x compared to naive implementation in CPU, Metal and NVIDIA GPU backends, respectively (Section  4).
@@ 2 Flash Memory & LLM Inference
In this section, we explore the characteristics of memory storage systems (e.g., flash, DRAM), and their implications for large language model (LLM) inference. We aim to understand the challenges and hardware-specific considerations essential for algorithm design, particularly in optimizing inference when working with flash memory.
@@ 2.2 Read Throughput
Flash memory systems perform optimally with large sequential reads. For instance, benchmarks on an Apple MacBook M1 Max with 1TB flash memory demonstrate speeds exceeding 6 GiB/s for a 1GiB linear read of an uncached file. However, this high bandwidth cannot be achieved for smaller, random reads due to the inherent multi-phase nature of these reads, encompassing the operating system, drivers, interrupt handling, and the flash controller, among others. Each phase introduces latency, disproportionately affecting smaller reads.
To circumvent these limitations, we advocate two primary strategies, which can be employed jointly. The first involves reading larger chunks of data. For smaller blocks, a substantial part of the overall read time is spent waiting for data transfer to begin. This is often referred to as latency to first byte. This latency reduces the overall throughput of each read operation considerably because the overall measured throughput has to take into account not just the speed of transfer once it begins, but the latency before it begins as well, which penalizes small reads. This means that if we coalesce the reads for rows and columns of the FFN matrices, we can pay the latency cost only once for any given row/column pair in both matrices and higher throughput can be realized. This principle is depicted in Figure  2(b). Perhaps a counterintuitive yet interesting observation is that in some scenarios, it will be worthwhile to read more than needed (but in larger chunks) and then discard, rather than only reading strictly the necessary parts but in smaller chunks. The second strategy leverages parallelized reads, utilizing the inherent parallelism within storage stacks and flash controllers. Our results indicate that throughputs appropriate for sparse LLM inference are achievable on modern hardware using 32KiB or larger random reads across multiple threads.
Motivated by the challenges described in this section, in Section  3, we propose methods to optimize data transfer volume and enhance read throughput to significantly enhance inference speeds.
@@
Figure 3: (a) Preactivations of tokens in one sequence in OPT 6.7B. The blue graph shows the preactivation of elements that the predictor detected as positive while the green graph is for up projection. As it can be seen most of the False Positives are close to 0 and False Negatives constitute a small portion of the elements. (b) A small low-rank predictor finds out which intermediate neurons are going to be activated.
@@ 3 Load From Flash
Our proposed solutions for reducing latency under memory constraints are categorized into areas:
  1. 1.
    Reducing Data Load: Aiming to decrease latency associated with flash I/O operations by loading less data [^1] It is notable that, by data we often refer to the weights of the neural network. However, the techniques we have developed can be easily generalized to other data types transferred and used for LLM inference, such as activations or KV cache, as suggested by Sheng et al. (2023)..
  2. 2.
@@
It is important to note that our focus is not on optimizing the compute, as it is orthogonal to the core concerns of our work. Instead, we concentrate on optimizing flash memory interactions and memory management to achieve efficient inference on memory-constrained devices. We will elaborate on the implementation details of these strategies in the experimental setup section.
@@ 3.1 Reducing Data Transfer
Our method leverages the inherent activation sparsity found in Feed-Forward Network (FFN) models, as documented in preceding research. The OPT 6.7B model, for instance, exhibits a notable 97% sparsity within its FFN layer. Similarly, the Falcon 7B model has been adapted through fine-tuning, which involves swapping their activation functions to ReLU, resulting in 95% sparsity while being similar in accuracy (Mirzadeh et al., 2023). Replacing activations of Llama 2 model (Touvron et al., 2023b) by FATReLU and finetuning can achieve 90% sparsity(Song et al., 2024). In light of this information, our approach involves the iterative transfer of only the essential, dynamic subset of the weights from flash memory to DRAM for processing during inference.
Selective Persistence Strategy. We opt to retain the embeddings and matrices within the attention mechanism of the transformer constantly in DRAM. For the Feed-Forward Network (FFN) portions, only the non-sparse segments are dynamically loaded into DRAM as needed. Keeping attention weights, which constitute approximately one-third of the model’s size, in memory, allows for more efficient computation and quicker access, thereby enhancing inference performance without the need for full model loading.
Anticipating ReLU Sparsity. The ReLU activation function naturally induces over 90% sparsity in the FFN’s intermediate outputs, which reduces the memory footprint for subsequent layers that utilize these sparse outputs. However, the preceding layer, namely the up project, must be fully present in memory.
To avoid loading the entire up projection matrix, we follow  Liu et al. (2023b), and employ a low-rank predictor to identify the elements zeroed by ReLU (see Figure  3(b)). We used a balanced loss over negative and positive samples of each layer. In contrast to their work, our predictor needs only the output of the current layer’s attention module and not the previous layer’s FFN module. We have observed that postponing the prediction to the current layer is sufficient for hardware-aware weight-loading algorithm design but leads to more accurate outcomes due to deferred inputs. We used 10000 samples from the C4 training dataset to do the training for 2 epochs. It took 4 hours on an A100 GPU to train each predictor.
We thereby only load elements indicated by the predictor, as shown in Figure  3(a). Furthermore, as demonstrated in Table  1, using predictors does not adversely affect the model’s performance in 0-shot tasks. For more details please refer to Appendix B.
Table 1: The low-rank predictor has a marginal impact on zero-shot metrics as the predictor of each layer accurately identifies sparsity.
@@
Figure 4: (a) Aggregated neuron usage of the tenth layer of Falcon 7B: the slope of aggregated neuron usage is decreasing. Other layers exhibit the same pattern. (b) Rather than deleting neurons that were brought to DRAM we keep the active neurons of past $k$ tokens (we use $k=5$): when the new token "Was" is being processed only a small fraction of new weights need to be loaded.
The Sliding Window Technique. In our study, we define an active neuron as one that yields a positive output in our low-rank predictor model. Our approach focuses on managing neuron data by employing a Sliding Window Technique. This technique entails maintaining a DRAM cache of only the weight rows that were predicted to be required by the recent subset of input tokens. The key aspect of this technique is the incremental loading of neuron data that differs between the current input token and its immediate predecessors. This strategy allows for efficient memory utilization, as it frees up memory resources previously allocated to cached weights required by tokens that are no longer within the sliding window (as depicted in Figure  4(b)).
From a mathematical standpoint, let $s_{\text{agg}}(k)$ denote the cumulative use of neuron data across a sequence of $k$ input tokens. Our memory architecture is designed to store an average of $s_{\text{agg}}(k)$ in DRAM. As we process each new token, the incremental neuron data, which is mathematically represented as $s_{\text{agg}}(k+1)-s_{\text{agg}}(k)$, is loaded from flash memory into DRAM. This practice is grounded in the observed trend of decreasing aggregated neuron usage over time. Consequently, larger values of $k$ result in a lesser volume of data being loaded for each new token (refer to Figure  4(a)), while smaller values of $k$ can help conserve DRAM that is used to store the cached weights. In determining the size of the sliding window, the aim is to maximize it within the constraints imposed by the available memory capacity.
3.2 Increasing Transfer Throughput
@@
Bundling Columns and Rows. Note that in the FFN layer, the usage of the $i$th column from the up projection and the $i$th row from the down projection coincides with the activation of the $i$th intermediate neuron. Consequently, by storing these corresponding columns and rows together in flash memory, we can consolidate the data into larger chunks for reading. Refer to Figure  5 for an illustration of this bundling approach. If each element of weights of the network is stored in num_bytes such bundling doubles the chunk size from $d_{model}\times$num_bytes to $2d_{model}\times$num_bytes as shown in Figure  5. Our analysis and experiment show this increases the throughput of the model.
Bundling Based on Co-activation. We hypothesized that neurons might exhibit highly correlated activity patterns, enabling bundling. By analyzing activations on the C4 validation dataset, we found a power law distribution of coactivations. However, bundling neurons with their highest coactivated neuron (closest friend) led to multiple loadings of highly active neurons, counteracting our goal. This result suggests that very active neurons are the closest friends of many others. We present this negative result to inspire future research on effective neuron bundling for efficient inference. Please refer to Appendix D for details.